Find the total differential of z=xsiny-ysinx


1
Expert's answer
2021-07-08T09:59:42-0400
"\dfrac{\partial z}{\partial x}=\sin y-y\cos x"

"\dfrac{\partial z}{\partial y}=x\cos y-\sin x"

"dz=(\sin y-y\cos x)dx+(x\cos y-\sin x)dy"